iger am (d) (4 points) Calculate the table of expected counts for these data. 15% (e) (2 points) Calculate the value of the x2 statistic for testing your null hypothesis from (a). You must show your work for full credit. 50. 1.8 (0 (I point) Using statistical software, compute the p-value associated with the x2 statistic from part (d). (g) (2 points) Based on the results to (c) and (d), what can you conclude about your hypothesis from part (a)? Give an explanation in the context of this situation.
Answer #1

d)

df=(row-1)(column-1) = (2-1)(2-1)=1

chi square stat = 7.8

excel formula for p-value

=CHISQ.DIST.RT(7.8,1)

=0.005224

e)

since pvalue is less tha alpha=0.05,so reject Ho

there is enough evidence to conclude that there is a relationship between anger and heart disease